40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Jan 28 18:03:53 2021, shruti, DailyProgress, PSOMA, cavity locking Path.pdf
    Reply  Mon Feb 1 18:25:47 2021, shruti, DailyProgress, PSOMA, cavity locking 
       Reply  Wed Feb 3 10:56:54 2021, shruti, Update, PSOMA, cavity locking 
          Reply  Thu Feb 4 16:19:35 2021, shruti, Update, PSOMA, cavity alignment 5FC37CEF-5B30-4FF0-ADF6-9FD4C59A1027.jpeg
             Reply  Thu Feb 4 22:27:43 2021, rana, Update, PSOMA, cavity alignment IMC_HOManalysis.ipynb
                Reply  Tue Feb 9 11:14:48 2021, shruti, Update, PSOMA, cavity alignment 
                   Reply  Tue Feb 16 11:30:14 2021, shruti, Update, PSOMA, Beat recovered, fiber organization, alignment F2E33DA5-DB03-46EC-B341-E388D96F1BC5.jpeg069D0D53-9111-4B33-96AD-05B907C4BBCB.jpegBeat_low_freq_oscillations.mp4
Message ID: 2623     Entry time: Mon Feb 1 18:25:47 2021     In reply to: 2620     Reply to this: 2626
Author: shruti 
Type: DailyProgress 
Category: PSOMA 
Subject: cavity locking 

Some Observations:

  • Temperature of laser (in kOhms) North: 10.056, South: 10.097
  • Power (in mW) North: 7.02, South: 7.26
  • Beat frequency: ~20 MHz, -50 dBm
  • After changing south laser temperature to 10.985 kOhms, the beat moved to 40 MHz, 4 dBm (!)

Some information:

  • Ring cavity made of :

Ma R~90%, Mb R>96%, Mc R>96% RoC 1m Coastline Optics CO-SAP-1000-250-1.0MP

Ma -> Mb -> Mc -> Ma

  • Photo-diode: PDA20CS

Updates:

  • Got the beam in the triangular cavity to more or less align, i.e., overlap with itself as checked on one side of the triangle
  • I hooked up the transmission PD in the transmission path of Mc
    • At different temperature settings, blocking the incoming beam (Mb -> Mc) resulted in the power becoming minimum and blocking the beam (Mc -> Ma) resulted in some variable fraction of the the nominal power which I take to be a sign of interference (?)
    • I moved around the temperature using the setpoint knob, but since the movement of the knob was nonlinear which along with a nonlinear change in power and frequency (with temperature) made it hard for me to figure out exactly what was happening in the limited time I spent on this so far

Next steps:

  • Going to explore the laser lock on the Moku:Lab for a while, after which I'll decide whether or not to continue with that or move to mixers + PDH servo box

 

ELOG V3.1.3-